What is the [H3O+ ] of 1.0 M  HF mixed with 1.0 M NaF. Ka = 6.6 x 10-4


1
Expert's answer
2022-05-06T13:50:04-0400

Solution:

The Henderson-Hasselbalch equation is commonly used to calculate the pH of a buffer solution from the concentration of the buffer components:



where pKa is the acid dissociation constant and [base] and [acid] are the base and acid concentrations, respectively.


base = NaF

acid = HF

Therefore,

pH = −log(6.6×10−4) + log(1.0 / 1.0) = 3.18 + 0 = 3.18

pH = 3.18


pH = −log[H3O+]

Therefore,

[H3O+] = 10−pH = 10−3.18 = 6.6×10−4

[H3O+] = 6.6×10−4 M


Answer: [H3O+] = 6.6×10−4 M